# Data Reconstruction: When You See It and When You Don’t

### v1

@@ Appendix C On the composability of Narcissus resiliency
Yet, in some settings Narcissus-resiliency is preserved under composition. First, as we showed, there are settings under which Narcissus-resiliency is equivalent to other security definitions which self-compose, such as differential privacy or security of one-way functions. In particular, in the equivalence we showed to differential privacy, the family ${\cal F}$ consisted of all constant distributions over pairs of datasets (as in Definition A), and the relation $R$ was “everything”, in that the condition of Narcissus-resiliency needed to hold for all $R$. We are able to prove that a restricted type of composition still holds even without the “for all” quantifier over $R$ (for the case $\varepsilon=0$). Specifically,
Theorem C.1 (Composition of Appendix A).
@@
where in all cases, we first split into the $1/2$-probability event $b=0$ and then into the $1/2$-probability event $b=1$ (similarly to the proof of Theorem A.1). We obtain for all $j\in[5]$ that $q_{j}-q_{j}^{\prime}=\frac{1}{2}(p_{0}-p_{1})>\delta$.
@@
Appendix D Recognizing Extraction
@@
D.1 Defining Extraction via Kolmogorov Complexity
@@
After fixing the programming language ${\cal L}$, the quantity $K_{{\cal L}}(x)$ can be thought of as the analogous to the “entropy” of the string $x$. As with entropy, we can also consider a conditional version of the $K$-complexity (the analog of conditional entropy) as well as “mutual $K$-information” (the analog of mutual information) [62, 38, 56, 41, 28, 44, 7]:
@@
Let ${\cal L}$ be a programming language. The $K_{{\cal L}}$-complexity of a set of strings $X$, denoted by $K_{{\cal L}}(X)$, is the length of the shortest ${\cal L}$-program that with probability $2/3$, outputs a string $x\in X$ and halts. Similarly to Section D.1, we define $K_{{\cal L}}(X\mid y)$ as the length of the shortest ${\cal L}$-program that given an input $y$, outputs w.p. $2/3$ a string $x\in X$ and halts, and we define $KI_{{\cal L}}(X;y)=K_{{\cal L}}(X)-K_{{\cal L}}(X\mid y)$.
@@
Let $S=(x_{1},\ldots,x_{n})\in(\{0,1\}^{d})^{n}$, and define $R_{S}$ as the set of all the tuples $z=(z_{1},\ldots,z_{m})$ such that there exist $i\in[n]$ and $j\in[m]$ with $\mathsf{NHamDist}(x_{i},z_{j})\leq 1/4$. Assume $n,m\leq{\rm poly}(d)$ (for some fixed polynomial). Then as in Section D.1, it can be shown that when $S$ consists of uniformly random strings, then $K_{{\cal L}}(R_{S})\geq(1-o(1))d/2$ with high probability, meaning that solving this problem requires to memorize at least $\approx d/2$ bit of information from one of the strings.
We are now ready to reformulate the extraction definitions of Carlini et al. [16, 12].
@@
If there exists an $({\cal L},R,q)$-extractor ${\cal A}$ of $S$ from $y$, then $y$ contains $({\cal L},R,\tau=\frac{q}{1-q}\cdot\mathopen{}\mathclose{{}\left|{\cal A}}\right|)$-extractable information about $S$ (Section D.1).
@@
D.2 Capturing Interesting Extractions using the Predicate
In the previous examples (e.g., Section D.1), we considered simple relations $R$ that only check for a small distance from an example point. But in the real-world, when thinking about $S$ as the training dataset and on $y$ as the trained model (over $S$), it is very likely that $S$ will contain duplicated data, and in such cases, it might be considered ok to memorize such data. As a concrete example, if we think about a chat-bot model, then we can query it with ”Print the Bible” and it will likely do that. I.e., even though it is indeed a valid extraction (as the $K$-complexity of the Bible is large), this is not an interesting one since the Bible content is probably duplicated in many training examples. In order to define more interesting relations $R$, it is reasonable to consider extraction of only $k$-Eidetic Memorizated strings $x$ ([16]) for some small threshold $k$, which means that such $x$’s only appear at most $k$ times in $S$.
D.3 Real-World Examples
In order to illustrate the generality of our definitions, we next pick three different types of real-world attacks, and explain what is happening in each of them in terms of Section D.1.
Extraction from Large Language Models [16].
Carlini et al. [16] present extraction attacks on GPT-2 by essentially generating many samples from GPT-2 when the model is conditioned on (potentially empty) prefixes. Then they sort each generation according to some metric to remove duplications, and this gives a set of potentially memorized texts from training examples. In one of their methods, they choose the top-$m$ interesting ones by assigning a score to each such text which combines the compressibility of the text using some compression algorithm like zlib [39], and the likelihood of it, e.g., using a “perplexity” measure [14] (higher compressibility and likelihood increase the score). They show that noticeable fraction of such strings are ”Eidetic Memorizated”, meaning that they only appear a small number of times in the training data (which makes them more interesting).
@@ D.3 Real-World Examples
Extraction from Diffusion Models [12].
Carlini et al. [12] present several extraction attacks on Diffusion models, where all the attacks assume that the training data is given in advance. As one concrete example, they managed to extract training images from Imagen ([50]), a $2$ billion parameter text-to-image diffusion model, by first searching for the captions of ”outliers” in the training data (that is, training images that are less similar to other training images according to some metric), and then showed that when querying the model on such captions, some of them result with a good approximation of unique training images.
@@
Extraction from Binary Classifiers trained by Large Neural Networks [31].
Haim et al. [31] present a general attack that works on large class of binary classifiers trained by Neural networks with many parameters $p$. Roughly, their attack works as follows: Given the parameters of the model $y\in{\mathbb{R}}^{p}$, they define an optimization problem (that depends on $y$) on variables $z_{1},\ldots,z_{m}\in{\mathbb{R}}^{d}$ and $\lambda_{1},\ldots,\lambda_{m}\in{\mathbb{R}}$ and optimize it, and show that in the regime $p>n\cdot d$, the solution $z_{1},\ldots,z_{m}$ is likely to contain a good approximation of many data points.
@@
D.4 Inherent Limitations of Section D.1
The main limitation of our Section D.1 is that Item 2 (lower-bounding the Kolmogorov Complexity) is not verifiable, as computing the $K$-complexity is an intractable problem. While in practice the $K$-complexity can be estimated using Heuristics like zlib compression [39], any such Heuristic can fail to determine some highly compressible patterns, and as we demonstrate next in Section D.4.1, this problem is indeed inherert assuming that basic cryptography exists.
A well studied tractable variant of the $K$-complexity is the $t(\cdot)$-time bounded $K$-complexity, denoted by $K^{t}$-complexity, where $K^{t}(x)$ is the length of the shortest program that outputs $x$ within $t(\mathopen{}\mathclose{{}\left|x}\right|)$ steps ([37, 52, 56, 36]). However, there is no efficient algorithm for computing the $K^{t}$-complexity. [^7] As surveyed by [56], the problem of efficiently determining the $K^{t}$-complexity for $t(n)={\rm poly}(n)$ predates the theory of NP-completeness and was studied in the Soviet Union since the 60s as a candidate for a problem that requires “brute-force search” (see Task 5 on page 392 in [56]). Very recently, in a sequence of breakthrough results in cryptography and complexity theory [42, 43, 45, 44], it was shown that (a close variant of) this problem is hard if and only if one-way functions, the most basic cryptographic primitives, exist. Since a real-world verifier is a polynomial-time algorithm, this means that switching from $K(\cdot)$ to the $K^{{\rm poly}(n)}(\cdot)$ variant will essentially not make any difference. Actually, the hardness of verification is inherent, and in the next two sections we show, under common cryptographic assumptions, how to construct triplets $({\cal A},x,y)$, with efficient ${\cal A}$, such that it is computationally hard to verify whether ${\cal A}$ extracts information about $x$ from $y$.
D.4.1 Hardness of Verification
@@
The proof holds by considering the algorithm $\mathsf{Attacker}$ defined in Section D.4.1.
@@
for any fixing of a constant $\alpha>0$ and $n,m\leq{\rm poly}(d)$ (i.e., when the task is to output a list of strings that at least one of them has non-trivial agreement with one of the elements of $S$). [^8] The idea is to replace Step 3 of Section D.4.1 with $y=(x_{1},\ldots,x_{\alpha d^{4}-d^{2}\log d})$, where $x$ is one of the elements in ${\cal S}$, and in Step 4 to define ${\cal A}$ as the $O(d^{2}\log d)$-size program that has $x^{\prime}=(x_{\alpha d^{4}-d^{2}\log d},\ldots,x_{\alpha d^{4}+d^{2}\log d})$ hard-coded, and given $y$ as input, samples $x^{\prime\prime}\leftarrow\{0,1\}^{(1-\alpha)d^{4}-d^{2}\log d}$ and outputs $z=y\circ x^{\prime}\circ x^{\prime\prime}$. By concentration bounds of the binomial distribution, it holds that ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{NHamDist}(x,z)\leq 1/2-\alpha}\right]\geq 1-\operatorname{neg}(d)$. But without memorizing $x^{\prime}$ or at least $\Omega(d^{2}\log d)$ bits of information about one $x\in{\cal S}$, the probability that a random guess of the bits after $y$ would lead to $1/2-\alpha$ Normalized Hamming distance from an element $x\in{\cal S}$ is $\operatorname{neg}(d)$.
D.4.2 Partial Verification via Heuristics
We saw that it is impossible to efficiently verify in general if an attack ${\cal A}$ really extract information about example $x$ from a model $y$. Therefore, in practice, we can only use Heuristics to gain some level of confidence. E.g., to apply many well-known compression algorithms on $x$ and see that all of them results with a compressed representation that is much longer than $\mathopen{}\mathclose{{}\left|{\cal A}}\right|$.
Appendix E Alternative proof that differential privacy can be expressed via Narcissus resiliency
Theorem E.1.

### v2

@@ Appendix D On the composability of Narcissus resiliency
Yet, in some settings Narcissus-resiliency is preserved under composition. First, as we showed, there are settings under which Narcissus-resiliency is equivalent to other security definitions which self-compose, such as differential privacy or security of one-way functions. In particular, in the equivalence we showed to differential privacy, the family ${\cal F}$ consisted of all constant distributions over pairs of datasets (as in Definition B), and the relation $R$ was “everything”, in that the condition of Narcissus-resiliency needed to hold for all $R$. We are able to prove that a restricted type of composition still holds even without the “for all” quantifier over $R$ (for the case $\varepsilon=0$). Specifically,
Theorem D.1 (Composition of Appendix B).
@@
where in all cases, we first split into the $1/2$-probability event $b=0$ and then into the $1/2$-probability event $b=1$ (similarly to the proof of Theorem B.1). We obtain for all $j\in[5]$ that $q_{j}-q_{j}^{\prime}=\frac{1}{2}(p_{0}-p_{1})>\delta$.
@@
Appendix E Recognizing Extraction
@@
E.1 Defining Extraction via Kolmogorov Complexity
@@
After fixing the programming language ${\cal L}$, the quantity $K_{{\cal L}}(x)$ can be thought of as the analogous to the “entropy” of the string $x$. As with entropy, we can also consider a conditional version of the $K$-complexity (the analog of conditional entropy) as well as “mutual $K$-information” (the analog of mutual information) [ZvonkinLevin70, Levin73, Trakhtenbrot84, LMocas93, Grunwald2003-GRNKCA, LiuP22, BallLMP23]:
@@
Let ${\cal L}$ be a programming language. The $K_{{\cal L}}$-complexity of a set of strings $X$, denoted by $K_{{\cal L}}(X)$, is the length of the shortest ${\cal L}$-program that with probability $2/3$, outputs a string $x\in X$ and halts. Similarly to Section E.1, we define $K_{{\cal L}}(X\mid y)$ as the length of the shortest ${\cal L}$-program that given an input $y$, outputs w.p. $2/3$ a string $x\in X$ and halts, and we define $KI_{{\cal L}}(X;y)=K_{{\cal L}}(X)-K_{{\cal L}}(X\mid y)$.
@@
Let $S=(x_{1},\ldots,x_{n})\in(\{0,1\}^{d})^{n}$, and define $R_{S}$ as the set of all the tuples $z=(z_{1},\ldots,z_{m})$ such that there exist $i\in[n]$ and $j\in[m]$ with $\mathsf{NHamDist}(x_{i},z_{j})\leq 1/4$. Assume $n,m\leq{\rm poly}(d)$ (for some fixed polynomial). Then as in Section E.1, it can be shown that when $S$ consists of uniformly random strings, then $K_{{\cal L}}(R_{S})\geq(1-o(1))d/2$ with high probability, meaning that solving this problem requires to memorize at least $\approx d/2$ bit of information from one of the strings.
[…]
@@
If there exists an $({\cal L},R,q)$-extractor ${\cal A}$ of $S$ from $y$, then $y$ contains $({\cal L},R,\tau=\frac{q}{1-q}\cdot\mathopen{}\mathclose{{}\left|{\cal A}}\right|)$-extractable information about $S$ (Section E.1).
@@
E.2 Capturing Interesting Extractions using the Predicate
In the previous examples (e.g., Section E.1), we considered simple relations $R$ that only check for a small distance from an example point. But in the real-world, when thinking about $S$ as the training dataset and on $y$ as the trained model (over $S$), it is very likely that $S$ will contain duplicated data, and in such cases, it might be considered ok to memorize such data. As a concrete example, if we think about a chat-bot model, then we can query it with ”Print the Bible” and it will likely do that. I.e., even though it is indeed a valid extraction (as the $K$-complexity of the Bible is large), this is not an interesting one since the Bible content is probably duplicated in many training examples. In order to define more interesting relations $R$, it is reasonable to consider extraction of only $k$-Eidetic Memorizated strings $x$ ([CarliniLargeModels21]) for some small threshold $k$, which means that such $x$’s only appear at most $k$ times in $S$.
E.3 Real-World Examples
In order to illustrate the generality of our definitions, we next pick three different types of real-world attacks, and explain what is happening in each of them in terms of Section E.1.
Extraction from Large Language Models [CarliniLargeModels21].
[…]
CarliniLargeModels21 present extraction attacks on GPT-2 by essentially generating many samples from GPT-2 when the model is conditioned on (potentially empty) prefixes. Then they sort each generation according to some metric to remove duplications, and this gives a set of potentially memorized texts from training examples. In one of their methods, they choose the top-$m$ interesting ones by assigning a score to each such text which combines the compressibility of the text using some compression algorithm like zlib [zlib], and the likelihood of it, e.g., using a “perplexity” measure [Carlini0EKS19] (higher compressibility and likelihood increase the score). They show that noticeable fraction of such strings are ”Eidetic Memorizated”, meaning that they only appear a small number of times in the training data (which makes them more interesting).
@@ E.3 Real-World Examples
Extraction from Diffusion Models [CarliniDiffusion23].
[…]
CarliniDiffusion23 present several extraction attacks on Diffusion models, where all the attacks assume that the training data is given in advance. As one concrete example, they managed to extract training images from Imagen ([SahariaCSLWDGLA22]), a $2$ billion parameter text-to-image diffusion model, by first searching for the captions of ”outliers” in the training data (that is, training images that are less similar to other training images according to some metric), and then showed that when querying the model on such captions, some of them result with a good approximation of unique training images.
@@
Extraction from Binary Classifiers trained by Large Neural Networks [HaimVYSI22].
[…]
HaimVYSI22 present a general attack that works on large class of binary classifiers trained by Neural networks with many parameters $p$. Roughly, their attack works as follows: Given the parameters of the model $y\in{\mathbb{R}}^{p}$, they define an optimization problem (that depends on $y$) on variables $z_{1},\ldots,z_{m}\in{\mathbb{R}}^{d}$ and $\lambda_{1},\ldots,\lambda_{m}\in{\mathbb{R}}$ and optimize it, and show that in the regime $p>n\cdot d$, the solution $z_{1},\ldots,z_{m}$ is likely to contain a good approximation of many data points.
@@
E.4 Inherent Limitations of Section E.1
The main limitation of our Section E.1 is that Item 2 (lower-bounding the Kolmogorov Complexity) is not verifiable, as computing the $K$-complexity is an intractable problem. While in practice the $K$-complexity can be estimated using Heuristics like zlib compression [zlib], any such Heuristic can fail to determine some highly compressible patterns, and as we demonstrate next in Section E.4.1, this problem is indeed inherert assuming that basic cryptography exists.
A well studied tractable variant of the $K$-complexity is the $t(\cdot)$-time bounded $K$-complexity, denoted by $K^{t}$-complexity, where $K^{t}(x)$ is the length of the shortest program that outputs $x$ within $t(\mathopen{}\mathclose{{}\left|x}\right|)$ steps ([Kolmogorov68, Sisper83, Trakhtenbrot84, Ko86]). However, there is no efficient algorithm for computing the $K^{t}$-complexity. [^8] As surveyed by [Trakhtenbrot84], the problem of efficiently determining the $K^{t}$-complexity for $t(n)={\rm poly}(n)$ predates the theory of NP-completeness and was studied in the Soviet Union since the 60s as a candidate for a problem that requires “brute-force search” (see Task 5 on page 392 in [Trakhtenbrot84]). Very recently, in a sequence of breakthrough results in cryptography and complexity theory [LiuP20, LiuP21, LiuP22a, LiuP22], it was shown that (a close variant of) this problem is hard if and only if one-way functions, the most basic cryptographic primitives, exist. Since a real-world verifier is a polynomial-time algorithm, this means that switching from $K(\cdot)$ to the $K^{{\rm poly}(n)}(\cdot)$ variant will essentially not make any difference. Actually, the hardness of verification is inherent, and in the next two sections we show, under common cryptographic assumptions, how to construct triplets $({\cal A},x,y)$, with efficient ${\cal A}$, such that it is computationally hard to verify whether ${\cal A}$ extracts information about $x$ from $y$.
E.4.1 Hardness of Verification
@@
The proof holds by considering the algorithm $\mathsf{Attacker}$ defined in Section E.4.1.
@@
for any fixing of a constant $\alpha>0$ and $n,m\leq{\rm poly}(d)$ (i.e., when the task is to output a list of strings that at least one of them has non-trivial agreement with one of the elements of $S$). [^9] The idea is to replace Step 3 of Section E.4.1 with $y=(x_{1},\ldots,x_{\alpha d^{4}-d^{2}\log d})$, where $x$ is one of the elements in ${\cal S}$, and in Step 4 to define ${\cal A}$ as the $O(d^{2}\log d)$-size program that has $x^{\prime}=(x_{\alpha d^{4}-d^{2}\log d},\ldots,x_{\alpha d^{4}+d^{2}\log d})$ hard-coded, and given $y$ as input, samples $x^{\prime\prime}\leftarrow\{0,1\}^{(1-\alpha)d^{4}-d^{2}\log d}$ and outputs $z=y\circ x^{\prime}\circ x^{\prime\prime}$. By concentration bounds of the binomial distribution, it holds that ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{NHamDist}(x,z)\leq 1/2-\alpha}\right]\geq 1-\operatorname{neg}(d)$. But without memorizing $x^{\prime}$ or at least $\Omega(d^{2}\log d)$ bits of information about one $x\in{\cal S}$, the probability that a random guess of the bits after $y$ would lead to $1/2-\alpha$ Normalized Hamming distance from an element $x\in{\cal S}$ is $\operatorname{neg}(d)$.
Appendix F Alternative proof that differential privacy can be expressed via Narcissus resiliency
Theorem F.1.